Some psychology students at Chaffey college wanted to know if there was a relationship between the number of selfies taken per day and narcissism. The narcissism scale was 1-6 Likert scale, where 4 is very narcissistic and 1 is not narcissistic at all. This was compared to the number of selfies taken per day.

You can use one of the charts in your packet or make your own.

The OBT is .681

Selfies

5

1

8

4

4

12

2

5

7

6

Narcissism

6

1

5

4

6

6

1

6

6

5

  • Make sure you also sketch a graph.
  • Do the appropriate Statistic.
  • Find Y for X=9, X=10, X =3.
Answer #1

X: Number of selfjer token per day Y: Narcissism X Y x2 XY 30 5 6 25 40 8 5 4 4 64 16 16 144 4 12. 6 6. 24 72 2 30 2 4 5 6 25When X = 9 Y=0.437 (9) + 2.24 0 Y = 6.173 When X=10. Y = 0.437 (10) +2.240 Y = 6.61 When X = 3 Y=0.437 (3) +2.240 Y = 3.8551